Assets:  $200 Reserves; $5000 Short term Bonds; $6000 Long Term Loans

Liabilities: $7000 Checkable Deposits; $3000 Fixed Rate Borrowings; $1200 Capital

If market interest rates fall by 2%, how will bank profits change?
Answer #1

Total interest-earning assets = Short term bonds + Long term loans = 5000 + 6000 = 11000

Total interest-paying liabilities = Checkable Deposits + Fixed rate borrowings = 7000 + 3000 = 10000

Since Total interest-earning assets > Total interest-paying liabilities, lower interest rate will decrease profit.

Decrease in profit = (Total interest-earning assets - Total interest-paying liabilities) x 2%

= (11000 - 10000) x 2%

= 1000 x 2%

= 20